# Yao Cheng, Qier Cui, Yunshi Lan, Xiang Li

### v1

@@ Abstract
Peer review, as a cornerstone of scientific research, ensures the integrity and quality of scholarly work by providing authors with objective feedback for refinement. However, in the traditional peer review process, authors often receive vague or insufficiently detailed feedback, which provides limited assistance and leads to a more time-consuming review cycle. If authors can identify some specific weaknesses in their paper, they can not only address the reviewer’s concerns but also improve their work. This raises the critical question of how to enhance authors’ comprehension of review comments. In this paper, we present SEAGraph, a novel framework developed to clarify review comments by uncovering the underlying intentions behind them. We construct two types of graphs for each paper: the semantic mind graph, which captures the author’s thought process, and the hierarchical background graph, which delineates the research domains related to the paper. A retrieval method is then designed to extract relevant content from both graphs, facilitating coherent explanations for the review comments. Extensive experiments show that SEAGraph excels in review comment understanding tasks, offering significant benefits to authors.
@@
Jianxiang Yuthanks:   Equal Contribution, Jiaqi Tan11footnotemark: 1, Zichen Ding, Jiapeng Zhu, Jiahao Li, Yao Cheng, Qier Cui, Yunshi Lan, Xiang Li East China Normal University, Shanghai, China sea.ecnu@gmail.com
@@ 1 Introduction
In recent years, the number of academic publications has grown exponentially, creating a vast “sea of papers” Bornmann and Mutz (2015); Lin et al. (2023). Traditionally, authors rely on the peer review process to receive feedback on their manuscripts Lee et al. (2013); Björk and Solomon (2013). However, the review cycle typically requires several months or even longer, which is time-consuming Horbach and Halffman (2018). Meanwhile, the large volume of submissions results in uncertain review qualities Bornmann and Mutz (2015), often resulting in ambiguous or overly brief comments that are challenging to explain Verma et al. (2022). For example, the statement that “the method is limited” is very vague without any details provided. In such cases, authors may be confused about which parts of the paper to be revised, leading to the difficulty in effectively addressing the reviewers’ issues or fully identifying the shortcomings of the paper. If they receive detailed suggestions, they may benefit more during this review process. Therefore, it is essential to understand the vague comments with more details provided, thereby helping the authors to improve their paper quality (as shown in Fig. 1).
@@
Currently, Large Language Models (LLMs) have shown powerful text comprehension and generation capabilities Achiam et al. (2023); Wei et al. (2022), offering new directions for revealing the underlying intentions behind each review comment. A straightforward approach is to provide LLMs with both the comment and the corresponding paper. Yet, it is usually difficult to feed an entire paper into LLMs for identifying key points, as review comments typically focus on specific aspects rather than the entire paper. Another alternative approach is using RAG (Retrieval-Augmented Generation) Cheng et al. (2024); Jiang et al. (2023), which enhances reasoning by retrieving the most relevant passages from lengthy texts based on the query. Nevertheless, the information retrieved by RAG tends to be fragmented, lacking clear logic Cao et al. (2024). In contrast, review comments are given based on the coherent logical structure formed when reviewers read the paper, which cannot be easily captured by fragmented segments. Recently, the success of GraphRAG Edge et al. (2024), which splits lengthy texts into discrete chunks and hierarchically connects them, has inspired new directions. Similarly, papers are inherently structurally organized with sections and subsections provided. Therefore, we can format papers as structured graphs, from which logical chains can be extracted to facilitate a deeper understanding of review comments.
In this paper, we propose SEAGraph, a novel framework designed to uncover the intentions behind paper reviews and enhance the understanding of review comments. We construct two distinct graphs for each reviewed paper: a semantic mind graph and a hierarchical background graph. Building upon the principles of the mind map D’Antoni and Zipp (2006), which employs visuospatial orientation to integrate information, we introduce the semantic mind graph to facilitate deeper semantic connections and organization of key knowledge points. In addition, the hierarchical background graph connects various related papers based on the themes of the paper, thereby simulating its research context. After the construction of the two graphs, we design a tailored retrieval method to extract the most relevant content from both graphs in response to each review comment. The extracted content is subsequently fed into LLMs to generate coherent and logical arguments that explain the reviewer’s comments. Overall, our contributions are summarized as follows:
@@ 2.1 Retrieval-Augmented Generation
RAG (Retrieval-Augmented Generation) improves the generation performance of LLMs by incorporating external knowledge Lewis et al. (2020). Initially, naive RAG approaches follow a process including indexing, retrieval, and generation Li et al. (2022). Advanced RAG frameworks focus on enhancing the retrieval quality during the pre-retrieval and post-retrieval phase like query rewriting, query expansion, and chunk reranking  Ma et al. (2023); Peng et al. (2024b); Zheng et al. (2023). Furthermore, some modular RAG approaches put forward new modules or pipelines to enhance the retrieval capability and alignment with task-specific requirements Yu et al. (2022); Shao et al. (2023). Despite these advancements, RAG faces challenges in handling query-focused summarization tasks when queries target entire text corpora Cao et al. (2024). GraphRAG emerges as an innovative solution to address this challenge Peng et al. (2024a). Edge et al. (2024) establish logical relationships between segments by connecting chunks or communities through a hierarchical structure. Wu et al. (2024a) and Sepasdar et al. (2024) construct specialized knowledge graphs, extending GraphRAG to the medical and soccer domains. In this work, we construct two logically connected graphs for the paper, leveraging the strengths of GraphRAG to better address the review comment understanding tasks.
@@ 2.2 Large Language Models in Peer Review
Recently, Large Language Models (LLMs) have made remarkable progress in text generation tasks Zhao et al. (2023); Ouyang et al. (2022); Luo et al. (2024), prompting researchers to explore new opportunities in the field of peer review  Li et al. (2024b); Checco et al. (2021). A significant focus has been placed on generating automated reviews to enhance the quality of academic papers Gao et al. (2024). For example, Liu and Shah (2023) and Liang et al. (2023) customize prompts to guide GPT-4 in generating scientific feedbacks. Yu et al. (2024a) and Wei et al. (2023) respectively employ supervised fine-tuning and continuous pretraining to align LLMs with human reviews. Expanding on this research, Jin et al. (2024) employ LLMs to simulate the entire review process, thereby revealing the impact of various factors on academic evaluation. Subsequently, Ye et al. (2024) highlight the risks associated with using LLMs for automated peer reviewing. Building on these insights, Yu et al. (2024b) delve into the challenge of distinguishing between peer reviews authored by humans and those generated by LLMs. However, both human reviews and generated reviews may suffer from issues such as ambiguity or brevity, causing confusion for authors. Our work aims to leverage LLMs to understand the intent of review comments, thereby assisting authors in polishing their papers.
@@ 3 SEAGraph
This section provides a detailed description of each module in SEAGraph, and the overall framework is in Figure 2.
3.1 Review Comments Understanding Task
@@
Our dataset consists of the PDF versions of academic papers and their corresponding review comments. We begin by utilizing Nougat Blecher et al. (2023) as the parser, a model built on the Visual Transformer architecture specifically tailored for extracting information from academic documents. Then, we construct the Semantic mind graph and the Hierarchical background graph for each reviewed paper, denoted as $\mathcal{G}_{S}(\mathcal{V}_{S},\mathcal{E}_{S})$ and $\mathcal{G}_{H}(\mathcal{V}_{H},\mathcal{E}_{H})$, where $\mathcal{V}$ and $\mathcal{E}$ represent nodes and edges, respectively.
@@
We input the entire review into LLM to extract individual comments under three textual sections: “Strengths, “Weaknesses”, and “Questions.” Then, all review comments are defined as a query set $\mathcal{Q}$, where $q\in\mathcal{Q}$ represents a single comment.
Our Goal.
Our goal is to retrieve subgraphs from two graphs $\mathcal{G}_{S}$ and $\mathcal{G}_{H}$, which represent evidence relevant to a review comment $q$. These pieces of evidence, along with the comment, are then provided to LLMs to help understand the details of the review. By generating a logical chain, SEAGraph enables authors to better understand reviewers’ perspectives and proceed with subsequent research more effectively.
@@ 3.2 Semantic Mind Graph Construction
We first use the Spacy library AI (2017) to break a full paper down to sentence level, allowing us to assess the relevance between sentences and decide whether adjacent sentences should be merged into chunks. Specifically, we utilize Sentence-BERT Reimers and Gurevych (2019) to encode both sentences and chunks, and design a semantic relevance measure to determine whether the current chunk is related to the next sentence. Subsequently, we place the first sentence $s_{1}$ into the initial chunk $c_{\text{current}}$. For each subsequent sentence $s_{i}$, we compute the embedding similarity between $s_{i}$ and $c_{\text{current}}$. If the similarity exceeds a threshold $\theta_{1}$, we merge the sentence into the current chunk by:
@@ 3.3 Hierarchical Background Graph Construction
Now, for a reviewed paper, we construct its hierarchical background knowledge graph based on theme nodes, abstract nodes, and semantic mind graphs. The first level includes multiple theme nodes, each corresponding to a thematic description that encapsulates the research topics. The second level connects these theme nodes to abstract nodes, where each abstract serves as a concise summary of a paper, representing its key ideas and maintaining a direct association with its respective theme. The third level extends from the abstract nodes to semantic mind graphs, which provide fine-grained information, offering a deeper insight into the paper’s structure and details. This hierarchical design clearly delineates the logical progression from themes to papers and further to detailed information, forming a systematic framework for representing the research background.

### v2

# Yao Cheng, Qier Cui, Yunshi Lan, Yao Liu, Xiang Li

@@ Abstract
Peer review, as a cornerstone of scientific research, ensures the integrity and quality of scholarly work by providing authors with objective feedback for refinement. However, in the traditional peer review process, authors often receive vague or insufficiently detailed feedback, which provides limited assistance and leads to a more time-consuming review cycle. If authors can identify some specific weaknesses in their paper, they can not only address the reviewer’s concerns but also improve their work. This raises the critical question of how to enhance authors’ comprehension of review comments. In this paper, we present SEAGraph, a novel framework developed to clarify review comments by uncovering the underlying intentions behind them. We construct two types of graphs for each paper: the semantic mind graph, which captures the authors’ thought process, and the hierarchical background graph, which delineates the research domains related to the paper. A retrieval method is then designed to extract relevant content from both graphs, facilitating coherent explanations for the review comments. Extensive experiments show that SEAGraph excels in review comment understanding tasks, offering significant benefits to authors. By bridging the gap between reviewers’ critiques and authors’ comprehension, SEAGraph contributes to a more efficient, transparent and collaborative scientific publishing ecosystem.
@@
Jianxiang Yuthanks:  Equal Contribution, Jiaqi Tan11footnotemark: 1, Zichen Ding, Jiapeng Zhu, Jiahao Li, Yao Cheng, Qier Cui, Yunshi Lan, Yao Liu, Xiang Lithanks:   Corresponding author East China Normal University, Shanghai, China sea.ecnu@gmail.com https://github.com/ecnu-sea/SEAGraph
@@ 1 Introduction
In recent years, the number of academic publications has grown exponentially, creating a vast “sea of papers” bornmann2015growth; lin2023moprd. Traditionally, authors rely on the peer review process to receive feedback on their manuscripts lee2013bias; bjork2013publishing. The peer review process typically lasts for several months or even longer horbach2018changing, yet the crucial rebuttal phase - the limited window for author-reviewer communication - remains disproportionately short. During this brief interaction period, authors and reviewers must navigate complex technical discussions through written exchanges alone, without the benefit of face-to-face dialogue that could facilitate mutual understanding verma2022lack. This time-constrained communication may leads to suboptimal outcomes: while reviewers may provide valuable insights, authors struggle to fully comprehend or effectively address these comments within the tight rebuttal timeframe. Therefore, when authors better understand reviewers’ perspectives, it not only leads to improved revisions but also makes the entire review process more productive and rewarding for both authors and reviewers (see Figure 1).
@@
Currently, Large Language Models (LLMs) have shown powerful text comprehension and generation capabilities achiam2023gpt; wei2022emergent, offering new directions for revealing the underlying intentions behind each review comment. A straightforward approach is to provide LLMs with both the comment and the corresponding paper. Yet, it is usually difficult to feed an entire paper into LLMs for identifying key points, as review comments typically focus on specific aspects rather than the entire paper. Another alternative approach is using RAG (Retrieval-Augmented Generation) cheng2024lift; jiang2023active, which enhances reasoning by retrieving the most relevant passages from lengthy texts based on the query. Nevertheless, the information retrieved by RAG tends to be fragmented, lacking clear logic cao2024lego. In contrast, review comments are given based on the coherent logical structure formed when reviewers read the paper, which cannot be easily captured by fragmented segments. Recently, the success of GraphRAG graphrag, which splits lengthy texts into discrete chunks and hierarchically connects them, has inspired new directions. Similarly, papers are inherently structurally organized with sections and subsections provided. Therefore, we can format papers as structured graphs, from which logical chains can be extracted to facilitate a deeper understanding of review comments.
In this paper, we propose SEAGraph, a novel framework designed to uncover the intentions behind paper reviews and enhance the understanding of review comments. We construct two distinct graphs for each reviewed paper: a semantic mind graph and a hierarchical background graph. Building upon the principles of the mind map DANTONI20062, which employs visuospatial orientation to integrate information, we introduce the semantic mind graph to facilitate deeper semantic connections and organization of key knowledge points. In addition, the hierarchical background graph connects various related papers based on the themes of the paper, thereby simulating its research context. After the construction of the two graphs, we design a tailored retrieval method to extract the most relevant content from both graphs in response to each review comment. The extracted content is subsequently fed into LLMs to generate coherent and logical arguments that explain the reviewer’s comments. Overall, our contributions are summarized as follows:
@@ 2.1 Retrieval-Augmented Generation
RAG (Retrieval-Augmented Generation) improves the generation performance of LLMs by incorporating external knowledge lewis2020retrieval. Initially, naive RAG approaches follow a process including indexing, retrieval, and generation li2022survey. The indexing phase involves cleaning and segmenting raw data into text chunks and encoding them into vector space. Retrieval and generation follow by encoding user queries, matching them with nearest chunks, and synthesizing context-aware responses using retrieved content. Advanced RAG frameworks focus on enhancing the retrieval quality during the pre-retrieval and post-retrieval phase like query rewriting, query expansion, and chunk reranking  ma2023query; peng2024large; zheng2023take. Furthermore, some modular RAG approaches put forward new modules or pipelines to enhance the retrieval capability and alignment with task-specific requirements yu2022generate; shao2023enhancing. Despite these advancements, RAG faces challenges in handling query-focused summarization tasks when queries target entire text corpora cao2024lego. GraphRAG emerges as an innovative solution to address this challenge peng2024graph. graphrag establish logical relationships between segments by connecting chunks or communities through a hierarchical structure. wu2024medical and sepasdar2024enhancing construct specialized knowledge graphs, extending GraphRAG to the medical and soccer domains. In this work, we construct two logically connected graphs for the paper, leveraging the strengths of GraphRAG to better address the review comment understanding tasks.
@@ 2.2 Large Language Models in Peer Review
Recently, LLMs have made remarkable progress in text generation tasks zhao2023survey; ouyang2022training; lbs3, prompting researchers to explore new opportunities in the field of peer review  li2024sentiment; checco2021ai. A significant focus has been placed on generating automated reviews to enhance the quality of academic papers reviewer2. For example, liu2023reviewergpt and liang2023can customize prompts to guide GPT-4 in generating scientific feedbacks, while yu2024automated and academicgpt refine LLMs through fine-tuning and pretraining. Expanding on this, jin2024agentreview simulate the review process with LLMs to analyze evaluation factors. Meanwhile, ye2024we highlight risks in automated peer review, and yu2024your explore challenges in distinguishing human and LLM-generated reviews. Besides, previous studies have highlighted various challenges in the peer review and rebuttal process. kuznetsov2024can note the time-intensive nature of peer review, requiring extensive discussions, while purkayastha2023exploring emphasize challenges in rebuttals due to language barriers and experience gaps. huang2023makes stress the need for rebuttals to address all reviewer concerns and reach consensus. These findings highlight the importance of improving authors’ understanding of feedback. Our work aims to leverage LLMs to understand the intent of review comments, thereby assisting authors in polishing their papers.
@@ 3 SEAGraph
Accurately simulating the perspective of a reviewer necessitates not only enabling LLMs to understand the content of the paper, but also equipping them with the background knowledge required for peer review. To this end, we design a Semantic Mind Graph and a Hierarchical Background Graph to model the two corresponding types of knowledge structures. The former captures the paper’s intrinsic arguments and evidence, while the latter formalizes the broader domain context essential for an informed critique. This explicit, dual-graph approach significantly aids LLMs in integrating these different knowledge types, thereby fostering a more robust and interpretable emulation of a reviewer’s reasoning. In the following, we present the details of each module in SEAGraph, with the overall framework illustrated in Figure 2.
3.1 Data Preprocessing
@@
Our dataset consists of the PDF versions of academic papers and their corresponding review comments. We begin by utilizing Nougat nougat as the parser, a model built on the Visual Transformer architecture specifically tailored for extracting information from academic documents. Then, we construct the Semantic mind graph and the Hierarchical background graph for each reviewed paper, denoted as $\mathcal{G}_{S}(\mathcal{V}_{S},\mathcal{E}_{S})$ and $\mathcal{G}_{H}(\mathcal{V}_{H},\mathcal{E}_{H})$, where $\mathcal{V}$ and $\mathcal{E}$ represent nodes and edges, respectively.
@@
We input the entire review into LLM to extract individual comments like “Strengths, “Weaknesses”, and “Questions.” Then, all review comments are defined as a query set $\mathcal{Q}$, where $q\in\mathcal{Q}$ represents a single comment.
@@ 3.2 Semantic Mind Graph Construction
We first use the Spacy library spacy2 to break a full paper down to sentence level, allowing us to assess the relevance between sentences and decide whether adjacent sentences should be merged into chunks. Specifically, we utilize Sentence-BERT sentence-bert to encode both sentences and chunks, and design a semantic relevance measure to determine whether the current chunk is related to the next sentence. Subsequently, we place the first sentence $s_{1}$ into the initial chunk $c_{\text{current}}$. For each subsequent sentence $s_{i}$, we compute the embedding similarity between $s_{i}$ and $c_{\text{current}}$. If the similarity exceeds a threshold $\theta_{1}$, we merge the sentence into the current chunk by:
@@ 3.3 Hierarchical Background Graph Construction
For a reviewed paper, we construct its hierarchical background knowledge graph based on theme nodes, abstract nodes, and semantic mind graphs. The first level includes multiple theme nodes, each corresponding to a thematic description that encapsulates the research topics. The second level connects these theme nodes to abstract nodes, where each abstract serves as a concise summary of a paper, representing its key ideas and maintaining a direct association with its respective theme. The third level extends from the abstract nodes to semantic mind graphs, which provide fine-grained information, offering a deeper insight into the paper’s structure and details. This hierarchical design clearly delineates the logical progression from themes to papers and further to detailed information, forming a systematic framework for representing the research background.